OpenStudy (stormoshawty):

The sum members of angle X and angle Y is 90. If the measure of angle X is 30 less than twice the measure of angle Y, what is the measure of angle X? a. 20 b. 35 c. 50 d. 65

OpenStudy (accidentalaichan):

2y - 30 = x

OpenStudy (accidentalaichan):

Establish equations x+y = 90 x = 2y - 30 ... Substitute the x = into the first equation 2y - 30 + y = 90 ... solve for y 3y - 30 = 90 3y = 90 + 30 3y = 120 y = 40 ... Substitute y into either equation and solve for x x = 2(40) - 30 = 80 - 30 = 50 ... Solution x = 50, y = 40 (50,40)

OpenStudy (accidentalaichan):

So your answer is C